# InfoLossQA: Characterizing and Recovering Information Loss in Text Simplification

### v1

@@ C.1 Model Recall of Human-written QA
1
How were the carbon nanoparticles administered to the participants? This study was a dose-escalation study, meaning that the carbon nanoparticles were administered in increasing doses sequentially to different groups of participants.
What type of study was conducted to test the effects of carbon nanoparticles? An investigator-initiated, randomized controlled, dose-escalation study was conducted.
Full
Disregard background explanation in the original.
2
What condition did the babies in the study have? The babies in the study were diagnosed with moderate-to-severe bronchiolitis.
How was the study conducted? The study was a double-blind, randomized controlled trial on infants (1 to 12 months) who were diagnosed in the emergency department with moderate-to-severe bronchiolitis.
Full
Fully recalled even if phrased differently
3
What statistical tests were used to analyze the results? The results were analyzed using a repeated measures (the procedures were conducted multiple times), mixed design ANOVA test (compares the results from three or more study groups).
What type of statistical analysis was used to analyze the results? A repeated measures, mixed design ANOVA was used to analyze the results.
Full
Disregard background explanation
4
What was done to prevent bias in the trial? The trial was designed as a double-blind trial, meaning neither the patients nor the researchers knew which group of participants had been assigned which combination of sedative drugs.
How was this study conducted? This study was conducted as a clinical trial at a single center. It was double-blinded, meaning that neither the researcher nor the participants knew which treatment each participant received, and it was randomized, meaning the participants were randomly assigned one of the treatments.
Full
Model QA is quite broad but includes the missing bit of information (double-blind).
5
What was the major criterion for patients to participate in the study? The 155 patients who participated in this study had a glaucoma-related diagnosis, meaning they had been diagnosed with an eye condition that causes vision loss.
What kind of health condition did all enrolled patients have? All enrolled participants were diagnosed with a glaucoma-related condition.
Full
Main subject is “major criterion” which the model QA identifies. 155 patients is not the subject of the question, and background explanation “vision loss” can be ignored.
6
How were the carbon nanoparticles administered to the participants? This study was a dose-escalation study, meaning that the carbon nanoparticles were administered in increasing doses sequentially to different groups of participants.
@@
Partial
The question asks for dose-escalation. The answer shows increasing doses, but it misses calling out “dose escalation.”
7
@@
What specific parts of the digestive system were observed? The study looked at the antrum (part of the upper stomach) and the second duodenal portion (lower part of the stomach).
In this study, where were the muscle contractions assessed? The muscle contractions were assessed in the antrum, which is a part of the stomach.
Partial
8
What other analyses were done of the rsFC strength? Correlation analyses were done to see if the rsFC strength under heroin exposure was connected to the rewarding effect and plasma (in the blood) concentrations of heroin.
@@
Did the researchers also look for a relationship between the strength of the brain connections and the level of morphine in the patients’ blood? Yes, the researchers studied whether there was a correlation between the strength of resting state functional connectivity (rsFC) and the plasma concentration of morphine, a major metabolite of heroin.
Partial
Rewarding effect is missing.
9
@@
What kind of tests were used to assess patients in this study? Participants were assessed with laboratory tests, the United Kingdom screening test […], the Michigan neuropathy screening score […], and the Michigan diabetic neuropathy score […]. In addition, […]
What method was used to assess the level of neuropathy in the patients? The Michigan neuropathy screening score was used to assess the level of neuropathy in all patients.
Partial
One model generates relatively narrow questions which can touch upon a small part of a larger reference question.
10
@@
What kind of factors is the study analyzing to determine how effective the fast track area (FTA) method is? The study is analyzing the impact of a fast tract area (FTA) using both effectiveness measures, such as waiting times and length of stay in the waiting room, and quality measures, such as mortality rates and LWBS rates (patients who leave without being seen by a medical professional).
What is the purpose of the fast track area? The purpose of the fast track area is to reduce waiting times, patient dissatisfaction, and morbidity by providing a separate area for less urgent cases, allowing for faster treatment and better patient flow.
No
Superficially a lot of overlap, but the answer here needs to make clear what is measured not what the fast track area can hypothetically do.
11
@@
What aspects of heart function were increased using the combination of LS and SNP drugs? Left ventricular systolic (the heart’s ability to pump blood) and diastolic (heart relaxation and how long it takes for the ventricles, two chambers of the heart, to fill with blood) functions including LVEF (how much blood is pumped out of the heart each time it contracts (beats)) improved using the combination of LS and SNP drugs.
What was the main conclusion of the study? The study shows that the LS and SNP combination really improved heart function during and after surgery. Using these drugs led to less need for other supporting drugs or equipment. Thus, we suggest using this drug combination before and during surgery.
No
Does not give an answer about the actual heart function.
12
@@
How much did the special questionnaire overestimate calcium intake compared to the 24-hour recall? The questionnaire overestimated the average total calcium intake by 221 mg/d (milligrams per day) compared to the 24-hour recall across racial groups.
What was the variability in daily calcium intake estimated by the 24-hour recall? The variability or standard deviation in daily calcium intake estimated based on the 24-hour dietary recall was 433 milligrams per day.
No
13
How did the other treatments perform in comparison to each other? The results showed that Group 2 that received only the laser treatment showed better results than Group 1 who received only a low-dose of potassium nitrate.
@@
What were the specifics of the treatment groups in the study? The treatment groups in the study were: Group 1 - 5% potassium nitrate, Group 2 - gallium-aluminum-arsenide diode laser with a wavelength of 980 nm, noncontact pulse mode, and power wattage of 0.5 W, Group 3 - combined 5% KNO3 and the diode laser, and Group 4 - placebo (control).
No
@@ C.2 Quality of Generated QA
Criterion 1: Question Givenness
@@
Criterion 2: Question Localization
@@
Criterion 3: Answer Simplicity
@@
Criterion 4: Answerability/Question Relevance
@@
Criterion 5: Answer Accuracy

### v2

@@ C.1 Model Recall of Human-written QA
1 How were the carbon nanoparticles administered to the participants? This study was a dose-escalation study, meaning that the carbon nanoparticles were administered in increasing doses sequentially to different groups of participants. What type of study was conducted to test the effects of carbon nanoparticles? An investigator-initiated, randomized controlled, dose-escalation study was conducted. Full Disregard background explanation in the original.
2 What condition did the babies in the study have? The babies in the study were diagnosed with moderate-to-severe bronchiolitis. How was the study conducted? The study was a double-blind, randomized controlled trial on infants (1 to 12 months) who were diagnosed in the emergency department with moderate-to-severe bronchiolitis. Full Fully recalled even if phrased differently
3 What statistical tests were used to analyze the results? The results were analyzed using a repeated measures (the procedures were conducted multiple times), mixed design ANOVA test (compares the results from three or more study groups). What type of statistical analysis was used to analyze the results? A repeated measures, mixed design ANOVA was used to analyze the results. Full Disregard background explanation
4 What was done to prevent bias in the trial? The trial was designed as a double-blind trial, meaning neither the patients nor the researchers knew which group of participants had been assigned which combination of sedative drugs. How was this study conducted? This study was conducted as a clinical trial at a single center. It was double-blinded, meaning that neither the researcher nor the participants knew which treatment each participant received, and it was randomized, meaning the participants were randomly assigned one of the treatments. Full Model QA is quite broad but includes the missing bit of information (double-blind).
5 What was the major criterion for patients to participate in the study? The 155 patients who participated in this study had a glaucoma-related diagnosis, meaning they had been diagnosed with an eye condition that causes vision loss. What kind of health condition did all enrolled patients have? All enrolled participants were diagnosed with a glaucoma-related condition. Full Main subject is “major criterion” which the model QA identifies. 155 patients is not the subject of the question, and background explanation “vision loss” can be ignored.
@@
7 What specific parts of the digestive system were observed? The study looked at the antrum (part of the upper stomach) and the second duodenal portion (lower part of the stomach). In this study, where were the muscle contractions assessed? The muscle contractions were assessed in the antrum, which is a part of the stomach. Partial
@@
8 What other analyses were done of the rsFC strength? Correlation analyses were done to see if the rsFC strength under heroin exposure was connected to the rewarding effect and plasma (in the blood) concentrations of heroin. Did the researchers also look for a relationship between the strength of the brain connections and the level of morphine in the patients’ blood? Yes, the researchers studied whether there was a correlation between the strength of resting state functional connectivity (rsFC) and the plasma concentration of morphine, a major metabolite of heroin. Partial Rewarding effect is missing.
@@
9 What kind of tests were used to assess patients in this study? Participants were assessed with laboratory tests, the United Kingdom screening test […], the Michigan neuropathy screening score […], and the Michigan diabetic neuropathy score […]. In addition, […] What method was used to assess the level of neuropathy in the patients? The Michigan neuropathy screening score was used to assess the level of neuropathy in all patients. Partial One model generates relatively narrow questions which can touch upon a small part of a larger reference question.
@@
10 What kind of factors is the study analyzing to determine how effective the fast track area (FTA) method is? The study is analyzing the impact of a fast tract area (FTA) using both effectiveness measures, such as waiting times and length of stay in the waiting room, and quality measures, such as mortality rates and LWBS rates (patients who leave without being seen by a medical professional). What is the purpose of the fast track area? The purpose of the fast track area is to reduce waiting times, patient dissatisfaction, and morbidity by providing a separate area for less urgent cases, allowing for faster treatment and better patient flow. No Superficially a lot of overlap, but the answer here needs to make clear what is measured not what the fast track area can hypothetically do.
@@
11 What aspects of heart function were increased using the combination of LS and SNP drugs? Left ventricular systolic (the heart’s ability to pump blood) and diastolic (heart relaxation and how long it takes for the ventricles, two chambers of the heart, to fill with blood) functions including LVEF (how much blood is pumped out of the heart each time it contracts (beats)) improved using the combination of LS and SNP drugs. What was the main conclusion of the study? The study shows that the LS and SNP combination really improved heart function during and after surgery. Using these drugs led to less need for other supporting drugs or equipment. Thus, we suggest using this drug combination before and during surgery. No Does not give an answer about the actual heart function.
@@
12 How much did the special questionnaire overestimate calcium intake compared to the 24-hour recall? The questionnaire overestimated the average total calcium intake by 221 mg/d (milligrams per day) compared to the 24-hour recall across racial groups. What was the variability in daily calcium intake estimated by the 24-hour recall? The variability or standard deviation in daily calcium intake estimated based on the 24-hour dietary recall was 433 milligrams per day. No
@@
13 How did the other treatments perform in comparison to each other? The results showed that Group 2 that received only the laser treatment showed better results than Group 1 who received only a low-dose of potassium nitrate. What were the specifics of the treatment groups in the study? The treatment groups in the study were: Group 1 - 5% potassium nitrate, Group 2 - gallium-aluminum-arsenide diode laser with a wavelength of 980 nm, noncontact pulse mode, and power wattage of 0.5 W, Group 3 - combined 5% KNO3 and the diode laser, and Group 4 - placebo (control). No
@@ C.2 Quality of Generated QA
Criterion 1: question givenness.
@@
Criterion 2: question localization.
@@
Criterion 3: answer simplicity.
@@
Criterion 4: answerability/question Relevance.
@@
Criterion 5: answer accuracy.
@@
Appendix D Data Release and License
We reused RCT abstracts from the Evidence Inference V2.0 dataset (DeYoung et al., 2020b; evidence-inference.ebm-nlp.com, accessed 2024-05-28). After personal communication with the authors, it was confirmed that all articles in this dataset were from the PubMed Open Access subset which only includes license terms that allow reuse (ncbi.nlm.nih.gov/pmc/tools/openftlist, accessed 2024-05-28). After discussion with our institutions’ librarian on fair use, we release the annotations in InfoLossQA under CC-BY-4.0.